# Task Aware Dreamer for Task Generalization in Reinforcement Learning

### v4

@@ 4.3 Theoretical Analyses
Algorithm Hypothesis Cartpole-balance Walker-stand&walk Cheetah-run&run__\__back
&balance__\__sparse &prostrate&flip &flip&flip__\__back
@@
Algorithms Hypothesis Cheetah__\__speed Pendulum__\__angle Walker__\__speed
@@
5 Experiments
We now present empirical results to answer the following questions:
  • How about TAD’s generalization ability when handling image-based and state-based observations? (Sec. 5.3)
  • Can TAD be extended to more general settings like dynamic generalization? (Sec. 5.4)
5.1 Experimental Setup
Image-based Control.
To verify our analyses of TDR in Sec. 4.3, we consider several task combinations in DeepMind Control suite (DMC) (Tassa et al. 2018): (1) Cartpole-balance&balance__\__sparse, which shares the same optimal actions with TDR 0; (2) two task combinations with non-zero TDR of Walker-stand&walk&prostrat&flip, and Cheetah-run&run__\__back&flip&flip__\__flip__\__back, which are widely used in multi-task unsupervised RL (Sekar et al. 2020; Laskin et al. 2021). For example, Cheetah-run and Cheetah-flip hope a two-leg robot to move forward by running and flip around the torso, respectively (see Fig. 3), yielding almost opposite optimal Q functions with huge TDR. In Appendix C.1, we introduce more details about these task combinations.
To evaluate generalization with image-based observations, we extend tasks in DMC and design three task distributions: (1) Cheetah__\__speed($\alpha,\ \beta$), which extends Cheetah-run and hopes the agent to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$; (2) Pendulum__\__angle($\alpha,\ \beta$), extending Pendulum-swingup to keep pendulum’s pole within the target angle interval $(\arccos\alpha,\ \arccos\beta)$; and (3) Walker__\__speed($\alpha,\ \beta$), which is based on Walker-run and requires the planar walker to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$. For each task distribution, we sample $4$ training tasks and $2$ additional test tasks. More details are in Appendix C.2.
State-based Control.
To demonstrate the scalability of TAD, we also consider some state-based continuous robotic control task distributions simulated via MuJoCo (Todorov, Erez, and Tassa 2012). Following previous work (Finn, Abbeel, and Levine 2017), we choose tasks distributions: (1) Half-Cheetah-Fwd-Back, which owns two opposite tasks; and (2) Half-Cheetah-Vel and Humanoid-Direc-2D, which are task distributions with 100 training tasks 30 test tasks. More details of these task distributions are in Appendix D.
Baselines.
In DMC, we choose two model-free methods employing Markovian policies $\Pi_{1}$: CURL (Laskin, Srinivas, and Abbeel 2020) and SAC+AE (Yarats et al. 2021). In addition, we choose two classic world models, PlaNet (Hafner et al. 2019b) and Dreamer (Hafner et al. 2019a), which utilize historical state-actions in policies and belong to $\Pi_{2}$. Moreover, we take a SOTA model-based meta RL method belonging to $\Pi_{3}$: MAMBA (Rimon et al. 2024). In state-based control, besides PlaNet and Dreamer, we take some meta RL methods like MAML (Finn, Abbeel, and Levine 2017), RL2 (Duan et al. 2016), and VariBAD (Zintgraf et al. 2019), including zero-shot and few-shot evaluation, as reference.
Metrics.
For the task combinations, we evaluate the average return of all tasks to verify TDR. For the task generalization settings, we train agents in training tasks and evaluate their generalization abilities in test tasks. For all experiments, we repeat 5 different random seeds and report the mean $\pm$ std to mitigate the effects of randomness following previous works (Hafner et al. 2019a; Rimon et al. 2024).
@@ 5.1 Experimental Setup
Figure 4: The t-SNE clustering of state embeddings for different tasks sampled via Dreamer, TAD-CE, and TAD-SC.
5.2 Experimental Results for TDR
To validate our analyses of TDR, we report different algorithms’ performance of different task combinations in Table 1. Results show that TAD-CE and TAD-SC outperform all baselines, especially in those environments with high TDR. As derived in Theorem 3, in task combinations where TDR is 0 like Cartpole-balance&balance__\__sparse, different tasks share the same optimal action and $\Pi_{1},\Pi_{2}$ own the optimal policy, thus baselines like CURL and Dreamer perform well. Conversely, in other task combinations, different tasks’ optimal Q functions may differ a lot (like Chetah-run and Cheetah-run__\__back) and TDR is significantly huge. Thus methods with policies in $\Pi_{1},\Pi_{2}$ like CURL and Dreamer can not differentiate different tasks and perform poorly. Aligning with Theorem 3, baselines utilizing $\Pi_{3}$ like MAMBA outperform other baselines and TAD improves performance conspicuously.
Moreover, we present some visualization to better understand how TAD works. Fig, 3 shows video predictions of TAD for different tasks (Cheetah-run and Cheetah-flip). We use the same agent trained by TAD to sample trajectories for these two tasks and show their trajectories in lines 1 and 3 respectively. Given the first 7 steps as the context (their observations are similar but the rewards are different), we directly imagine the future 55 steps by our trained reward-informed world model, of which the results are in lines 2 and 4 of Fig. 3 respectively. As shown here, TAD performs differently to handle different tasks only by receiving different rewards. Moreover, our reward-informed world models demonstrate long-term prediction and high-quality reconstruction capabilities, which reveals the potential for training future large-scale world models. In Fig. 4, we sample trajectories from agents trained by Dreamer, TAD-CE, as well as TAD-SC, and visualize the states of different tasks, of which the dimensions are reduced for visualization by t-SNE (Van der Maaten and Hinton 2008). As shown here, states in TAD-CE and TAD-SC of different tasks are clearly distinguished, while Dreamer can not differentiate them and perform the same in different tasks. This result demonstrates TAD can effectively learn task-aware information and distinguish different tasks. Videos of different tasks are provided in supplementary materials.
@@ 5.3 Experimental Results for Task Generalization
To answer the second question, we report the task generalization results of image-based and state-based environments in Table 2-3. In Table 2, TAD shows more powerful generalization abilities than all baselines. This demonstrates that TAD can both handle multiple training tasks simultaneously and generalize to unseen test tasks effectively. We also provide some visualization results in Appendix C.3 with videos in supplementary materials.
Additionally, in Table 3, we compare TAD with existing meta RL methods in state-based environments, of which the training timesteps are 1e7, 1e7, and 1e6, respectively. Many context-based meta RL methods, including RL2, PEARL, and VariBAD, utilize historical rewards and belong to $\Pi_{3}$, thus they can distinguish different tasks, which also verify our analyses. As shown in Table 3, TAD achieves a significant improvement for both training tasks and test tasks, since TAD is aware of task information for generalizing to unseen tasks.
@@ 5.4 Ablation Study
In Fig. 5, we do ablation studies about the task term in TAD, i.e., we consider Dreamer(w/ r) that only integrates $\Pi_{3}$ into Dreamer. Results show that just providing rewards can help Dreamer distinguish different tasks, which verifies our analyses in TDR. Also, TAD-CE and TAD-SC show superior performance, demonstrating that our task model and corresponding ELBO are effective for task generalization. Also, we do ablation studies on reward signals for model-free methods like CURL in Appendix E.1.
Figure 5: Ablation study on Reward Signal.
Extension to Dynamic Generalization.
To answer the third question, we evaluate TAD in more general settings with different observations, dynamics, and/or actions. As TAD utilizes all historical information to infer the environment, it can be directly applied to these settings. We design task distribution with different embodiments (Acrobot-Cartpole-Pendulum, Walker-Cheetah-Hopper), as well as different dynamics (Cheetah-run__\__mass, Walker-walk__\__mass). More details about environments and results are in Appendix E.2, where TAD achieves much greater performance compared to baselines and show potential in further handling dynamic generalization and even cross-embodiment tasks.
@@ 5.5 Limitations and Discussion
In terms of limitations, TAD assumes that the task context is continuously related to historical information for generalizing to unseen tasks. Thus TAD might be difficult to generalize in sparse-reward settings. We further demonstrate that without extra knowledge or finetuning, zero-shot generalization to unseen tasks with extremely sparse rewards is impossible since there is no way to distinguish different tasks (Appendix A.7). Fortunately, in relatively sparse reward settings, we conduct experiments in Appendix E.3 to show that TAD can infer the current task and generalize to unseen tasks well.
@@ 6 Conclusion
In this work, we propose a novel framework of TAD that handles different tasks via all historical information and utilizes novel reward-informed world models to capture invariant latent features. In TAD, we calculate the corresponding variational lower bound of the data log-likelihood, which includes a novel loss term to distinguish different tasks via states. To explain components in TAD, we introduce a novel metric TDR to capture the relevance of the task distribution and show that Markovian policies perform poorly in tasks with high TDR. Experiments in image-based and state-based settings demonstrate that TAD can remarkably improve the performance of handling different tasks meanwhile, especially for high TDR ones, and successfully generalize to unseen tasks.
@@ A.1 The Proof of Theorem 1
Proof.
First, we will show that $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$, following the proof in (Young et al. 2023) that considers the setting that the rewards of all environments are the same.
$\begin{split}&\{q_{m}\}\in\mathcal{H}_{2}\\ \Leftrightarrow&\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall o,a\\ \Rightarrow&\exists(r_{m})_{m=1}^{M}::p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o_{t}^{m},a_{t}^{m}),a^{\prime}),\forall m,t\\ \Rightarrow&q_{m}(o_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(o_{t+1}^{m},a^{\prime}),\forall m,t\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{1},\end{split}$ (10)
@@
$\begin{split}&\{q_{m}\}\in\mathcal{H}_{3}\\ \Leftrightarrow&\exists p:p(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\forall m,t\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p(o,a),a^{\prime}),\forall m,o,a\\ \Rightarrow&\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\forall m,t\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall m,o,a\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{2},\end{split}$ (11)
thus we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$. ∎
@@ A.2 The Derivation of the ELBO
We use ${\color[rgb]{0,139,69}q}$ to represent $q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})$, ${\color[rgb]{0,0,1}\hat{q}}$ to present $q(s_{t}|o_{\leq t},a_{<t},r_{<t})$, ${\color[rgb]{1,0,0}\tilde{q}}$ to represent $q(s_{t-1}|o_{\leq(t-1)},r_{<(t-1)},a_{<(t-1)})$, and we have
$\begin{split}&\ln p(o_{1:T},r_{1:T},\mathcal{M}|a_{1:T})\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T},\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T}|\mathcal{M},s_{1:T})p(\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\right]\\ =&\ln\mathbb{E}_{{\color[rgb]{0,139,69}q}}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\frac{p(s_{t}|s_{t-1},a_{t-1})}{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}\right]\\ \geq&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})+\sum_{t=1}^{T}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ +&\mathbb{E}_{{\color[rgb]{0,139,69}q}}\sum_{t=1}^{T}\left[\ln p(s_{t}|s_{t-1},a_{t-1})-\ln q(s_{t}|o_{\leq t},r_{<t},a_{<t})\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]\\ +&\sum_{t=1}^{T}\left[\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})-\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ -&\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}{\color[rgb]{1,0,0}\tilde{q}}}\left[\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ -&\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{1,0,0}\tilde{q}}}[\mathrm{KL}\left(q(s_{t}|o_{\leq t},r_{<t},a_{<t})\|p(s_{t}|s_{t-1},a_{t-1})\right)].\end{split}$ (12)
Thus we have proven the ELBO.

### v5

@@ 5.1 Experimental setup
Algorithm Hypothesis Cartpole-balance Walker-stand&walk Cheetah-run&run_\_back
&balance_\_sparse &prostrate&flip &flip&flip_\_back
@@
Algorithms Hypothesis Cheetah_\_speed Pendulum_\_angle Walker_\_speed
@@
5.2 Experimental results for TDR
To validate our analysis of TDR, we report the performance of different algorithms across different task combinations in Table 1. The results show that TAD-CE and TAD-SC outperform all baselines, particularly in environments with high TDR. As illustrated by Theorem 4.4, in task combinations with zero TDR, since different tasks share the same optimal action, policy hypothesis sets $\Pi_{1},\Pi_{2}$ are sufficient to obtain the optimal policy. Therefore, baselines like CURL and Dreamer also perform well in such scenarios, e.g., Cartpole-balance&balance_\_sparse. Conversely, in task combinations where the optimal Q-functions vary significantly across tasks, such as Cheetah-run and Cheetah-run_\_back, the TDR is significantly huge. In these cases, methods with policies in $\Pi_{1},\Pi_{2}$ (e.g., CURL and Dreamer) cannot differentiate between tasks and perform poorly. Consistent with Theorem 4.4, baselines leveraging $\Pi_{3}$ (e.g., MAMBA) outperform other baselines, and TAD further improves performance substantially.
Moreover, we provide qualitative visualizations to further illustrate how TAD operates. Fig. 2 presents future predictions generated by TAD for two tasks (Cheetah-run and Cheetah-flip). Using the same agent trained with TAD, we collect trajectories for both tasks and display them in the first and third rows, respectively. Given the first 7 steps as context, where the observations are visually similar while the rewards differ, we roll out the next 55 steps purely in imagination using the trained RIWM. The corresponding predicted trajectories are shown in the second and fourth rows of Figure 2. As observed, TAD produces distinct predictions across tasks solely based on the reward signals, indicating that the agent successfully infers task context from historical information. In addition, the accurate long-horizon predictions and high-quality reconstructions produced by RIWM highlight its potential as a foundation for training larger-scale world models.
In Figure 3a, we further visualize the latent states of world models trained with Dreamer, TAD-CE, and TAD-SC. Trajectories from different tasks are embedded into a two-dimensional space using t-SNE [52]. The latent representations from TAD-CE and TAD-SC form clearly separated clusters across tasks, whereas Dreamer produces overlapping embeddings and fails to distinguish task context. This visualization confirms that TAD effectively learns task-aware representations and differentiates between tasks at the representation level.
@@ 5.3 Experimental results for task generalization
To address the second question, we report task generalization results on both image-based and state-based environments in Table 23. In Table 2, TAD exhibits stronger generalization capability than all baselines, indicating that it can simultaneously handle multiple training tasks and effectively generalize to unseen test tasks. Additional qualitative visualizations are provided in Appendix C.3.
Furthermore, Table 3 compares TAD with existing meta-RL methods in state-based environments, where the training horizons are $1e7$, $1e7$, and $1e6$, respectively. Many context-based meta-RL methods, such as RL2, PEARL, and VariBAD, utilize historical rewards and thus belong to $\Pi_{3}$. As a result, they are capable of distinguishing different tasks, which is consistent with our theoretical analysis. As shown in Table 3, TAD achieves substantial performance gains on both training and test tasks, demonstrating that leveraging task-aware representations enables more effective generalization to unseen tasks.
@@ 5.4 Ablation study
In Fig. 3b, we present ablation studies on the task-context term in TAD. In particular, we evaluate Dreamer(w/ r), which augments Dreamer to $\Pi_{3}$ by incorporating reward conditioning. The results show that simply providing reward signals can help Dreamer distinguish different tasks to some degree, which is consistent with our analysis in Sec. 4.3. Nevertheless, TAD-CE and TAD-SC achieve notably stronger performance, indicating that the proposed RIWM and the corresponding optimization method further enhance task generalization ability.
We also conduct ablations on reward signals for model-free methods such as CURL and SAC+AE, both of which belong to $\Pi_{1}$. Specifically, we design CURL (w/ r) and SAC+AE (w/ r) by directly appending the reward to the observation at each timestep. As shown in Table 1, their performance remains comparable to the original CURL and SAC+AE implementation. This is because these variants only use the reward and observation from the current timestep, without leveraging historical information, and thus still fail to effectively distinguish between different tasks.
(a) T-SNE visualization.
(b) Ablation study on Reward Signal.
Figure 3: (a) The t-SNE clustering of state embeddings for different tasks via Dreamer, TAD-CE, and TAD-SC. (b) The ablation study on the reward signals.
Extension to dynamics generalization.
To address the third question, we evaluate TAD in more general settings involving variations in observations, dynamics, and/or action spaces. Since TAD infers task information with all historical information, it can be directly applied to these settings without modification. We construct task distributions with different embodiments (Acrobot-Cartpole-Pendulum, Walker-Cheetah-Hopper), as well as different dynamics (Cheetah-run_\_mass, Walker-walk_\_mass). More details and results are provided in Appendix E.2, where TAD demonstrates substantially stronger performance than the baselines, highlighting its potential for dynamics generalization and even cross-embodiment adaptation.
@@ 5.5 Limitations and discussion
Regarding limitations, TAD assumes that task context can be inferred from historical interaction data in order to generalize to unseen tasks. Consequently, its effectiveness may be reduced in environments with extremely sparse rewards. In Appendix A.7, we further show that without extra prior knowledge or fine-tuning, zero-shot generalization to unseen tasks with extremely sparse rewards is fundamentally infeasible, as the agent lacks sufficient information to distinguish among tasks. Nevertheless, in relatively sparse-reward settings, our experiments in Appendix E.3 demonstrate that TAD can still infer the current task and generalize effectively to unseen tasks.
@@ 6 Conclusion
In this work, we propose Task-Aware Dreamer (TAD), a novel framework that leverages full historical information to handle multiple tasks and employs Reward-Informed World Models (RIWM) to capture task-specific latent features. In TAD, we derive a variational lower bound on the data log-likelihood that incorporates a task context term, encouraging the world model to learn task-specific representations. To justify the necessity of TAD’s components, we first introduce the Task Distribution Relevance (TDR) metric to characterize the relevance of different tasks from the task distribution. We then theoretically show that Markovian policies perform poorly in task distributions with high TDR. Extensive experiments in both image-based and state-based settings demonstrate that TAD substantially improves performance in the task distribution setting, particularly for task distributions with high TDR, and successfully generalizes to unseen tasks in a zero-shot manner.
[…]
This work was supported by the NSFC Projects (Nos. 92370124, 92248303, U25B6003, 62350080), Beijing Natural Science Foundation L247011, Tsinghua Institute for Guo Qiang, and the High Performance Computing Center, Tsinghua University.
@@ A.1 The proof of Theorem 1
Proof A.1.
First, we will show that $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$, following the proof in [39] that considers the setting where the rewards of all environments are the same.
$\begin{split}\{q_{m}\}\in\mathcal{H}_{2}\Leftrightarrow&\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall o,a\\ \Rightarrow&\exists(r_{m})_{m=1}^{M}::p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},q_{m}(o_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o_{t}^{m},a_{t}^{m}),a^{\prime}),\forall m,t\\ \Rightarrow&q_{m}(o_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(o_{t+1}^{m},a^{\prime}),\forall m,t\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{1},\end{split}$ (11)
@@
$\begin{split}\{q_{m}\}\in\mathcal{H}_{3}\Leftrightarrow&\exists p:p(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\forall m,t\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p(o,a),a^{\prime}),\forall m,o,a\\ \Rightarrow&\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\forall m,t\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall m,o,a\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{2},\end{split}$ (12)
thus we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$.
@@ A.2 The derivation of the ELBO
We use $\color[rgb]{0,139,69}q$} to represent $(s_{1:T}|o_{1:T},a_{1:T},r_{1:T}), q^{\hat{q}}$ to present $(s_{t}|o_{\leq t},a_{<t},r_{<t}), q~{\color[rgb]{1,0,0}\tilde{q}}$ to represent $q(s_{t-1}|o_{\leq(t-1)},r_{<(t-1)},a_{<(t-1)})$, and we have
$\begin{split}&\ln p(o_{1:T},r_{1:T},\mathcal{M}|a_{1:T})\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T},\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T}|\mathcal{M},s_{1:T})p(\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\right]\\ =&\ln\mathbb{E}_{{\color[rgb]{0,139,69}q}}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\frac{p(s_{t}|s_{t-1},a_{t-1})}{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}\right]\\ \geq&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})+\sum_{t=1}^{T}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]+\mathbb{E}_{{\color[rgb]{0,139,69}q}}\sum_{t=1}^{T}\left[\ln p(s_{t}|s_{t-1},a_{t-1})-\ln q(s_{t}|o_{\leq t},r_{<t},a_{<t})\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\left[\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})-\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]-\sum_{t=1}^{T}\mathbb{E}_{{\hat{q}}{\color[rgb]{1,0,0}\tilde{q}}}\left[\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]-\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{1,0,0}\tilde{q}}}[\mathrm{KL}\left(q(s_{t}|o_{\leq t},r_{<t},a_{<t})\|p(s_{t}|s_{t-1},a_{t-1})\right)].\end{split}$ (13)
Thus, we have proven the ELBO.